OpenStudy (anonymous):

In the 1996 Olympics, Michael Johnson of the United States won the gold medal in the 200-meter race with a time of 19.32 seconds. In the 2008 Olympics, Usain Bolt of Jamaica, won the gold medal in the 200-meter race in 19.30 seconds. If they ran the same race repeating their respective times, calculate the distance, in meters, that Bolt would have beat Johnson in this race? Its just .207 right?

OpenStudy (whpalmer4):

MJ goes 200m in 19.32 seconds, or 200m/19.32 seconds = 10.352 m/s UB goes 200m in 19.30 seconds. MJ in 19.30 seconds goes 200m/19.32 seconds * 19.30 seconds = 199.793 m so he has 200 m - 199.793 m = 0.207 m left to go when UB breaks the tape.